OpenStudy (anonymous):

Find the solution set for the following problem. 5 diminished by 3 times a number is at most 11. {x | x ≤ -2} {x | x ≥ -2} {x | x ≤ 2} {x | x ≥ 2}

OpenStudy (dumbcow):

hmm looks like 5-3x <= 11 -6 <= 3x -2 <= x

OpenStudy (anonymous):

so the second one?

OpenStudy (apoorvk):

Let the number be 'x'. 5 diminished by by 3 times it means 5 reduced by 3x. and it's atmost 11. that means less than or equal to 11. So, 5-3x <=11 Solve this inequality, and find the right option. Post to confirm~! ;)